Paula invited 15 friends to his sweet 16 birthday party. She decided to use 16 small square tables to form a large Square so that she and her 15 friends would all have a place to sit. at the last minute for cousins arrive from out of town. How could Paula arrange the 16 tables into a new rectangular shape ( with no holes in the middle ) so that all 20 people will have a place in there will be no empty seats?
Answer by Shin123(626)

15+1=16. If it is a square, 16 people can sit around it. 8*2=16 tables and 8+2+8+2=10+10=20 people. You need to arrange the tables 8 by 2 so everyone can fit.